OpenStudy (samigupta8):

the roots of x^3+6x^2-12x+32=0 were reduced by h to obtain a new cubic equation with no second degree term then h=?

hartnn (hartnn):

let the root be 'x' it was reduced by 'h' so, put \(\huge x-h\) in place of 'x' in your equation what do u get ?

OpenStudy (samigupta8):

(x-h)^3+6(x-h)^2-12(x-h)+32=0

hartnn (hartnn):

now the 2nd degree term is not present in this so, put the co-efficient of x^2 =0

OpenStudy (samigupta8):

i m getting h=-2

hartnn (hartnn):

how ? -3hx^2+6x^2 =0

OpenStudy (samigupta8):

thek hai smjh aa gya mujhe glti kr di thi sign ki sorry

hartnn (hartnn):

no prob :)

OpenStudy (samigupta8):

thanks....once again

hartnn (hartnn):

welcome ^_^